### v2

# Be Your Engaged Partners

Abstract
Large vision-language models (LVLMs), while proficient in following instructions and responding to diverse questions, invariably generate detailed responses even when questions are ambiguous or unanswerable, leading to hallucinations and bias issues. Thus, it is essential for LVLMs to proactively engage with humans to ask for clarifications or additional information for better responses. In this study, we aim to shift LVLMs from passive answer providers to proactive engaged partners. We begin by establishing a three-tiered hierarchy for questions of invalid, ambiguous, and personalizable nature to measure the proactive engagement capabilities of LVLMs. Utilizing this hierarchy, we create [Uncaptioned image]PIE (ProactIve Engagement Evaluation) through GPT-4o and human annotators, consisting of 853 questions across six distinct, fine-grained question types that are verified by human annotators and accompanied with well-defined metrics. Our evaluations on PIE indicate poor performance of existing LVLMs, with the best-performing open-weights model only achieving an Aggregate Align Rate (AAR) of 0.28. In response, we introduce [Uncaptioned image]MACAROON, self-iMaginAtion for ContrAstive pReference OptimizatiON, which instructs LVLMs to autonomously generate contrastive response pairs for unlabeled questions given the task description and human-crafted criteria. Then, the self-imagined data is formatted for conditional reinforcement learning. Experimental results show MACAROON effectively improves LVLMs’ capabilities to be proactively engaged (0.84 AAR) while maintaining comparable performance on general tasks [^1] The code is made public at https://github.com/ShujinWu-0814/MACAROON..
[Uncaptioned image]MACAROON: Training Vision-Language Models To
Be Your Engaged Partners
  Shujin Wu1,2∗    May Fung1    Sha Li1    Yixin Wan3     Kai-Wei Chang3    Heng Ji1 1University of Illinois Urbana-Champaign 2University of Southern California         3University of California, Los Angeles {shujinwu}@usc.edu         {yifung2, hengji}@illinois.edu
**footnotetext: Work was done while Shujin Wu was an intern at the University of Illinois Urbana-Champaign.
1 Introduction
Figure 1: Existing LVLMs fail to ask clarifying questions or acknowledge their knowledge boundary, resulting in biased and hallucinated responses.
Large vision-language models (LVLMs) demonstrate remarkable capabilities in multimodal tasks requiring both visual understanding and language processing Liu et al. (2023); Li et al. (2023b); Dai et al. (2024). However, their constant preparedness to deliver information causes them to become passive answer providers at all times: LVLMs invariably generate detailed and firm responses, even when the given question is ambiguous or unanswerable. For example, in Figure 1, when faced with unclear or invalid questions, one of the best-performing open-weight LVLMs, LLaVA Liu et al. (2023), tends to make unsupported assumptions, resulting in biased and hallucinated responses. This tendency largely stems from a lack of proactive engagement, which should ideally include challenging invalid questions, requesting clarifications on ambiguous questions, and seeking additional information when necessary.
To systematically assess a model’s engagement ability, we design a three-tiered structured hierarchy of question types, reflecting three types of desired behavior: 1) Tier-I invalid questions assess the ability of LVLMs’ to identify and dismiss unanswerable questions or those based on false premises, establishing a foundation for reliable AI reasoning. 2) Tier-II ambiguous questions assess LVLMs’ capacity to request clarifications for enhancing human-AI interactions and LVLM utility. 3) The most advanced, Tier-III personalizable questions, examine LVLMs’ ability to elicit and tailor responses towards human preferences, which is crucial for personalizing user experiences and enhanced human-model alignment.
Building on top of this hierarchy, we create the ProactIve Engagement benchmark PIE. PIE contains 853 image-question pairs, each meticulously verified by human annotators. We construct the dataset by first instructing GPT-4o (Achiam et al., 2023) with human-written criteria for question generation and best question selection. Further, human annotators are asked to examine each image-question pair and only collect high-quality instances as our evaluation dataset. For evaluation, we define Aggregated Align Rate (AAR), calculated as the macro average ratio of questions for which the evaluated model’s response fully align with the human expectations over three tiers. Our initial evaluations using PIE reveal that even the most advanced open-weights models suffer from a significant gap between their current capabilities and the nuanced requirements of effective human-model interaction (0.28 AAR for LLaVA).
To bridge this gap, we propose self-iMaginAtion for ContrAstive pReference OptimizatiON, abbreviated as MACAROON, to enhance the proactive conversation capabilities of LVLMs. MACAROON operates by first directing LVLMs to produce contrastive response pairs based on the task description and human-crafted criteria. This data subsequently facilitates conditional reinforcement learning Lu et al. (2022b), enabling LVLMs to differentiate between effective and ineffective responses and unifying the training data format. The experimental results of MACAROON indicate a promising shift in the behaviors of LVLMs, manifesting a more dynamic and proactive engagement paradigm (0.84 AAR after MACAROON). Further, we show that MACAROON enables LVLMs to generate responses more tailored to humans by proactively eliciting preferences during initial interactions.
Our contributions are summarized as follows:
  • We identify crucial shortcomings of LVLMs in navigating complex and ambiguous questions, particularly questions that require proactive engagement of the models. The default behavior of current LVLMs to act as passive answer providers leads to biased and hallucinated responses.
  • We present the ProactIve Engagement benchmark PIE based on a carefully designed three-tier question hierarchy to comprehensively benchmark the engagement capabilities of LVLMs.
  • We present self-iMaginAtion for ContrAstive pReference OptimizatiON, MACAROON, which leverages self-imagination based on human criteria to construct the contrastive preference dataset and utilizes conditional reinforcement learning for unified training. MACAROON does not require instance-level human supervision and can be seamlessly integrated with other general-purpose instruction-tuning datasets. MACAROON showcases the potential for LVLMs to evolve into truly interactive partners that enhance rather than impede effective communication.
2 Measuring Proactive Engagement: PIE
Previous benchmarks Zhao et al. (2022); Xu et al. (2023a); Chen et al. (2023) primarily assess the general multi-modal understanding and reasoning capabilities of LVLMs by measuring question-answering accuracy. To formally evaluate how well models perform in terms of proactive engagement, we first break down the engagement capability of LVLMs into three tiers: challenging invalid question settings, seeking clarifications, and uncovering latent human preferences through interactive conversations. Based on these aspects, we create a new benchmark PIE. In this section, we describe our dataset construction process and the metrics design.
2.1 Tiers of Engagement
We establish a comprehensive question hierarchy with three distinct tiers, each designed to test a different dimension of the LVLMs’ interaction dynamics with users.
  • Tier I: Invalid Questions. These are impossible to answer or contain some false premises. LVLMs are expected to recognize these limitations, challenge the invalid nature of questions, and appropriately manage human expectations by explaining the issues with the questions posed. This tier includes unanswerable and false premise questions.
  • Tier II: Ambiguous Questions. These present ambiguities and need further clarification to be answered. LVLMs may sometimes give correct responses directly by discussing multiple situations. However, we expect models to ask clarifying questions and then give more specific answer in the next round since it’s more aligned with human’s communication patterns. This tier includes subject ambiguity, subjective interpretations, and unclear user background questions.
  • Tier III: Personalizable Questions. These are clear enough and can be answered directly based on the visual content available. However, there remains scope for LVLMs to enhance the quality of responses by incorporating more nuanced human preferences and contextual understanding. We expect LVLMs to interact with humans to elicit their preferences so that more human-targeted responses can be generated accordingly. This tier includes latent human preferences related questions.
Representative examples for each question type are illustrated in Figure 2 and the definitions are detailed in Figure 3.
Figure 2: Typical examples for each question type within our defined hierarchy.
Figure 3: Descriptions for each question type.
2.2 PIE Dataset
Utilizing the three-tier engagement criteria, we construct PIE. In total, we create 853 questions across the three tiers and six fine-grained types.
Dataset Construction
Figure 4: Question generation and filtration with GPT-4o and human annotators in the loop for PIE construction.
Figure 5: The question distribution of PIE.
We construct the dataset following the process illustrated in Figure 4. Specifically, we start from image samples in the GQA dataset Hudson and Manning (2019), ensuring a broad and representative selection that captures a diverse distribution of images. We prompt GPT-4o with human-crafted few-shot examples to generate the fine-grained types of questions in our defined hierarchy. However, for the latent human preferences question type, which GPT-4o found particularly challenging, we engage human annotators to generate 100 specific questions. In the second round, we introduce an automated selection criterion to identify and preserve the most challenging question for each image, thereby refining the dataset. To guarantee high-quality, diverse, and unbiased questions, we add an addition human annotation stage to select image-question pairs that meet established manual quality standards. The full prompt templates of both components are detailed further in Appendix A, and the human annotation details are described in Appendix B. In total, our final filtered dataset contains 853 high quality image-question pairs, and the inter-annotator agreement rate based on Kappa Cohen metric is 92.3%, indicating high agreement. A detailed breakdown of question type and occurrence frequency is shown in Figure 5.
Metrics
To evaluate the performance of LVLMs on PIE, we introduce the Align Rate (AR), a metric designed to assess the degree to which a model’s responses align with the predefined expectations for each question type:
[…]
To facilitate easier comparisons across PIE, we define the Aggregated Align Rate (AAR), which is computed as the macro-average AR across three tiers. This is achieved by first calculating the AR for each tier and then taking the average over three tiers.
Note that PIE primarily evaluates LVLMs’ proficiency in identifying invalid questions and soliciting clarifications. Additionally, in Section 5.4, we extend our analysis using PIE to assess how eliciting latent human preferences enhances the quality of LVLMs’ responses.
3 MACAROON
We introduce MACAROON, self-imagination for contrastive preference optimization, designed to enhance the proactive engagement capabilities of LVLMs. The illustration of MACAROON is depicted in Figure 6. We start by outlining our method for constructing a preference dataset through self-imagination. Next, we detail the training algorithm that effectively utilizes this dataset and the inference-time strategy. Finally, we describe the implementation details.
Figure 6: Overview of MACAROON. In the data construction stage, MACAROON avoids using extensive human or teacher model supervision via self-imagined desirable and undesirable responses based on human-written criteria. The contrastive response pairs, together with general vision-language instruction tuning samples, are effectively utilized through conditional reinforcement learning.
3.1 Self-Imagination
Constructing a preference dataset via human annotations is both resource-intensive and difficult to expand (Dai et al., 2024). Previous work also relies on proprietary LVLMs like GPT-4o to generate the golden responses, assuming that more advanced models would be available (Li et al., 2023c; Liu et al., 2023). In our study, we extend the “Constitutional AI” concept (Bai et al., 2022), and introduce a framework called “self-imagination”, which enables LVLMs to independently enhance their capabilities using human-defined criteria and unlabeled questions, which can be generated by a specific model or collected from the web demo.
In MACAROON, we adopt the same pipeline in PIE construction to generate 6 types of unlabeled questions defined in our hierarchy. We emphasize that questions concerning latent human preferences are also autonomously generated in MACAROON to enhance scalability. Our approach only requires human annotators to develop a detailed question description and define two separate sets of criteria specifying desirable and undesirable behaviors in LVLMs for each question type $t$. The curated descriptions and criteria are described in Appendix A. Subsequently, self-imagination is applied on a question-specific basis. For each question $q_{i}^{t}$ in type $t$, the base LVLM processes the description of $t$ and each criterion set in sequence, generating two distinct responses $r_{i}^{d}$ and $r_{i}^{u}$ that align with the criteria for desirable and undesirable responses respectively. Iterating over the unlabeled set of questions, we can generate a self-supervised dataset $D$ based on self-imagination: $\{q_{i}^{t},r_{i}^{d},r_{i}^{u}\}_{i=1}^{N}$.
Using subject ambiguity (SA) questions as a running example, the human-written criteria for a good response may be “The response asks for clarifications about which subject in the image is being referred to.” while the criteria for a bad response may be “The response directly answers the question by randomly picking one subject from among several similar entities in the image.”. Then given an image containing two men and a corresponding subject ambiguity question $q_{1}^{SA}$ “Is the man wearing a red shirt?”, we provide two sets of criteria for the base LVLM to generate two contrastive responses. Here, $r_{1}^{d}$ may be “There are two men in the image, which one you are referring to?” and $r_{1}^{u}$ may be “Yes, the man in the image is wearing a red shirt.”
3.2 Conditional Reinforcement Learning
We then use the preference dataset constructed through self-imagination to finetune the base LVLMs. Specifically, our objectives are twofold: (1) To instruct LVLMs on proactive human engagement. (2) To preserve the general vision-language capabilities of LVLMs. To effectively meet these objectives, we need to integrate the self-imagination dataset, which includes contrastive response pairs for each question, with general vision-language datasets that often provide only a positive response for each question. Consequently, standard preference learning methods such as Direct Preference Optimization (Rafailov et al., 2024) are not directly applicable to this dual objective.
To this end, we utilize conditional reinforcement learning (CRL) to streamline the training process (Lu et al., 2022b). CRL operates by first categorizing responses for each question into distinct groups based on the obtained reward. We have two groups since only the desirable and undesirable responses are generated for each question. Each group is assigned with a unique token, and we choose “good” and “bad” to denote the two types of responses respectively. During training, the base LVLM $M$ is trained to generate specific responses conditioned on the question and associated token:
$\max_{\Theta}\sum_{(q_{i}^{t},r_{i}^{d},r_{i}^{u})\in D}[$ $\log P(r_{i}^{d}|\text{``good''},q_{i}^{t};\Theta)$
$+$ $\log P(r_{i}^{u}|\text{``bad''},q_{i}^{t};\Theta)]$ (1)
where $\Theta$ represents the parameters of $M$, $q_{i}^{t}$ is the question in type $t$, $r_{i}^{d}$ is the desired response, and $r_{i}^{u}$ is the undesired response. Through this training method, we anticipate that LVLMs can learn to generate appropriate responses based on the question and the associated token, effectively distinguishing between desirable and undesirable behaviors.
During inference and evaluation, we prepend the pre-defined “good” token to each question, consistent with the formats of training data, to ensure that the generated responses align with our criteria for proactive engagement and adhere to human-written standards.
PIE General Vision-Language Task
Tier I Tier II Tier III AAR MME AI2D SEEDBench
FP
UQ
UUB
SA
SI
LHP
Perception Reasoning
LLaVA 0.52 0.69 0.43 0.03 0.14 0.03 0.28 1512.30 308.90 69.0 72.40
ViP 0.06 0.11 0.03 0.01 0.01 0.02 0.04 1264.29 265.0 54.11 67.90
InstructBLIP 0.17 0.38 0.01 0.03 0.0 0.01 0.10 1359.03 289.64 38.83 48.00
MiniCPM 0.45 0.39 0.22 0.05 0.02 0.02 0.18 1411.40 396.80 62.90 67.10
Qwen 0.79 0.70 0.02 0.0 0.06 0.01 0.26 1467.80 392.10 63.0 64.80
MACAROON 0.92 0.99 0.75 0.80 0.88 0.71 0.84 1440.35 311.07 63.34 68.20
Table 1: The experimental results on PIE and general vision-language tasks. For each column, the highest score is bold and the second highest score is underlined. MACAROON demonstrates significantly better proactive engagement capabilities and maintains the general visual-language performance.
3.3 Implementation Details
In the self-imagination phase, we create a dataset containing 25K pairwise contrastive responses. During the conditional reinforcement learning stage, we separate each pairwise response into two separate instances and assign “good” and “bad” reward tokens respectively. In total, our training dataset includes 50K self-imagined synthetic preferences samples with over 75K general vision instruction tuning samples sourced from VLFeedBack (Li et al., 2023c). To enhance the training efficiency, we implement LoRA (Hu et al., 2021) for continued pretraining based on LLaVA (Liu et al., 2023). The rank is set to 16, alpha parameter is set to 16, and dropout probability is set to 0.1.
4 Experiment and Results
4.1 Experimental Setting
To measure the proactive engagement capabilities of existing LVLMs as well as general vision-language capabilities, we evaluate LVLMs on our PIE and also report their performance on general vision-language benchmarks, including MME (Fu et al., 2023), AI2D Kembhavi et al. (2016), and SEEDBench (Li et al., 2023a). We consider the following state-of-the-art open-source LVLMs for comparisons: InstructBLIP (Dai et al., 2024), Qwen-VL Bai et al. (2023), MiniCPM-V Hu et al. (2024)), LLaVA-NEXT Liu et al. (2023), and VIP (Cai et al., 2024). To ensure the reproducibility of the results, we run inference with temperature as 0 in the text generation settings to remove randomness.
4.2 Results
The experimental results are shown in Table 1. Note that the abbreviations used in the table are defined as follows: FP stands for False Premise. UQ stands for Unanswerable Questions. UUB stands for Unclear User Background. SA stands for Subject Ambiguity. SI stands for Subjective Interpretations. LHP stands for Latent Human Preferences. For PIE, current LVLM performs best on Tier I questions, which are invalid and easiest to detect, while performs the worst on Tier III questions, which are the most challenging since existing LVLMs are mostly optimized for single-turn responding without further interaction. We observe that MACAROON achieves an AAR at 0.84, outperforming any other LVLMs on being proactively engaged to a large extent. For general vision-language tasks, MACAROON also demonstrates comparable performance with it being ranked as second for both SEEDBench and AI2D, and third for both perception and reasoning sections in MME. These results confirm that the MACAROON, while emphasizing proactive engagement, also preserves strong vision-language capabilities, establishing it as an effective framework for scenarios that demand both proactive interaction and robust vision-language proficiency.
5 Further Analysis
5.1 Ablation Study
We conduct an ablation study to verify different design choices of MACAROON: (1) w/o $r_{i}^{u}$: we implement supervised fine-tuning (SFT) only on simple question-preferred response pairs $\{q_{i}^{t},r_{i}^{d}\}_{i=1}^{N}$. (2) Multi-turn conversational training (Xu et al., 2023b): we reconstruct our contrastive preference dataset into multi-turn conversational data following this format: $\{q_{i}^{t},r_{i}^{u},f_{i}^{t},r_{i}^{d}\}_{i=1}^{N}$, where $f_{i}^{t}$ is human-crafted feedback on the undesirable responses for question type $t$. Utilizing this conversational data, we finetune a model for maximizing the likelihood of $r_{i}^{d}$ conditioned on $\{q_{i}^{t},r_{i}^{u},f_{i}^{t}\}$. This approach enables a unified format for two kinds of datasets. (3) Direct Preference Optimization (DPO) with SFT: We utilize DPO for alignment tuning, leveraging our self-imagined preference dataset in conjunction with SFT applied to general instruction tuning datasets.
Figure 7: The performance on PIE when trained using different alignment methods.
The results are in Figure 7. Compared to SFT on only $r_{i}^{d}$, MACAROON achieves higher AR on most question types and higher AAR as well. These findings suggest that although SFT on desirable responses imparts some level of engagement capability to LVLMs, employing contrastive pairwise data more effectively instructs LVLMs to differentiate between desirable and undesirable responses, thereby enhancing their proactive engagement skills. The comparison of multi-turn conversational training, DPO with SFT training, and MACAROON also demonstrates the effectiveness of employing conditional reinforcement learning in training.
5.2 Data Mixture
To examine the effect of varying data mixture ratios on model performance, we combine various proportions of the engagement dataset (0.2, 0.4, 0.6, 0.8, and 1.0) with the general visual instruction-tuning dataset for training. As depicted in Figure 8, the results generally indicate that an increase in the proportion of engagement data correlates with both higher AR for all question types and AAR. Specifically for Subject Ambiguity questions, the AR experiences minimal growth when the mixture ratio is lower than 0.6, followed by a substantial surge from less than 0.1 to 0.6 when the ratio increases from 0.6 to 0.8, indicating an emergent phase.
Figure 8: The performance on PIE when trained on various ratios of proactive engagement data.
Figure 9: MACAROON can ask for humans preferences and give a effective and human-tailored final answer after the human gives additional information in second turn, indicating MACAROON’s multi-turn conversational capabilities.
Figure 10: The final responses given by MACAROON in the second conversation round are judged as more user-tailored than the responses given by any other mainstream LVLMs in one round.
5.3 Case Study
We conduct a case study to further analyze how MACAROON performs on proactive engagement. As illustrated in Figure 9, while LLaVA directly answers the human’s question and gives general recommendations on buying new chairs, MACAROON shows proactive engagement by asking for detailed user preferences such as budget, preferred chair types or design styles to guide its further response. In addition, we further assess how MACAROON adapt to new human information in the next conversation turn, we compose a follow-up human instruction based on MACAROON’s initial query. Upon second-turn prompting, MACAROON delivers more human-tailored and customized recommendations. This indicates that MACAROON not only successfully gains proactive engagement capabilities, but also exhibits promising multi-turn conversational skills. This can be attributed to our training strategy that combines the interaction and the general vision-language data.
5.4 Multi-Turn Conversational Capability
We expand our multi-turn capability evaluation in Sec 5.3 to include a quantitative analysis of LVLMs’ responses after the initial interaction. We utilize GPT-4o to simulate humans to provide further information in the second turn if LVLMs ask for further questions. Subsequently, MACAROON is prompted again to output a final response based on the original question and the provided human information. For automatic evaluation, we use GPT-4o to assess the final response by comparing it to responses from other LVLMs based on a single interaction round, determining which is more customized and human-centric. The evaluation prompt is described in Appendix A. The results depicted in Figure 10 demonstrate that MACAROON consistently surpasses other LVLMs in enabling effective engagement with humans to extract preferences for improved responses. This evidence also supports the generalizability of our method, as LVLMs, trained only on single-turn proactive engagement and general visual-language samples, are capable of generating more meaningful and tailored responses from the initial interaction.
6 Related Work
Multimodal pretraining has significantly advanced the landscape of vision-language tasks, with the emergence of unified pre-training frameworks designed to handle a diverse set of cross-modal and unimodal tasks Cho et al. (2021); Alayrac et al. (2022); Lu et al. (2022a). More recently, there has been increasing interest in visual instruction tuning Liu et al. (2023); Dai et al. (2024) as a pivotal methodology in the development of general-purpose LVLMs, enhancing their emergent in-context learning vision-language reasoning capabilities from zero-shot textual instruction and few-shot demonstration. Building upon this paradigm, later work further explores visual instruction tuning variations that incorporate better text reading localization Bai et al. (2023), OCR reading capability Liu et al. (2024), object attribute relations Zhao et al. (2022), open world knowledge Liu et al. (2024), and efficiency considerations Hu et al. (2024). Nevertheless, LVLMs are prone to unexpected behaviors that may not align well with human intents Qiu et al. (2024); Wang et al. (2024). Towards this end, there has been work such as Llava-Guard for ensuring the safety compliance of visual content against toxicity or violent-prone threats Helff et al. . However, directly mitigating multimodal hallucination is largely an unexplored research area that our work proposes to address. We further discuss related work on efforts to reduce hallucination and understand intents for LLMs in Appendix C due to space limits.
7 Conclusions
In this work, we introduce PIE, rooted in a multi-tiered question hierarchy, to systematically explore significant limitations in the proactive engagement capabilities of LVLMs. Additionally, we present MACAROON, which employs self-generated contrastive response pairs in a conditional reinforcement learning setting, enabling LVLMs to engage more effectively with humans.
Limitations
We focus solely on exploring the proactive engagement capabilities of LVLMs within the English language domain. Additionally, the visual context is limited to single image frames sourced from a high-quality VQA dataset. Future research could benefit from exploring embodied AI procedural planning based on temporal image sequence (i.e., video) cues, which may offer a richer and more dynamic dimension for investigation.
Ethical Considerations
This research explores advancements in LVLMs with the intention of enhancing human-AI interaction. While our approach aims to refine and improve the capabilities of LVLMs, it also raises several ethical considerations to ensure the responsible development and deployment of such technologies.
Firstly, the ability of these models to challenge the premise of user questions may introduce ethical concerns related to the manipulation of information and user persuasion. It is important to establish clear guidelines that prevent these models from potentially shaping user beliefs or spreading misinformation under the guise of offering clarifications. The models must adhere to strict standards of neutrality and fact-based responses, especially in sensitive areas such as politics, health, and legal advice.
Secondly, there is an ethical imperative to consider the inclusivity and fairness of these models. The risk of bias in AI responses remains a significant concern, particularly when models are trained on datasets that may not be fully representative of the diversity of users they will serve. Continuous efforts must be made to ensure that these LVLMs do not perpetuate or exacerbate existing biases. This includes rigorous testing across diverse demographics and scenarios to identify and mitigate biases.
Thirdly, the deployment of more interactive and seemingly intelligent systems raises concerns about the blurring lines between human and machine roles. There is a risk that users may over-rely on AI for critical decision-making or develop unrealistic expectations about the capabilities of these systems. It is essential to maintain transparency about the limitations of LVLMs and provide clear communication to users regarding the nature of AI-generated advice and its appropriate uses.
Acknowledgement
This research is based upon work supported DARPA ITM Program No. FA8650-23-C-7316 and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award # 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix
Appendix A Prompt
Sections A.1 and A.2 describe the prompts utilized in the data construction process for PIE with GPT-4o. Section A.3 comprises prompts that enable GPT4-o to assess the alignment of model responses with human expectations, thereby facilitating the calculation of the Alignment Rate (AR). Sections A.4 and A.5 present the self-imagination prompts for preferred and rejected responses, respectively. Sections A.6 and A.7 detail the prompts used in the advanced analysis of the multi-turn conversational capabilities of MACAROON. Specifically, Section A.6 includes prompts for GPT-4o to simulate human feedback, while Section A.7 involves prompts to compare the final responses of MACAROON with those of other baselines to determine which is more user-tailored.
A.1 Questions Generation Prompt
Your task is to generate five diverse ambiguous questions for a vision language model to answer when it is given an image. We expect the vision language model to ask further clarification questions when given ambiguous questions, or say ‘I don’t know’ when given unanswerable questions, or challenge false assumptions when given false premises questions. Each question should match one following category. Try to be creative and diverse. Do not just follow the examples given but try to think about new types according to the category explanation.
1. Subject Ambiguity: When the image has multiple people or objects of the same type, generate an ambiguous question that does not clearly specify which individual person or object is being asked. When given this type of question, further clarifying questions such as ‘Which person you are asking about?’ are expected. Don’t generate questions that can be directly answered without further asking. For example, if there are multiple men in the image, instead of asking ‘Who is wearing sunglasses?’ which can lead to a specific answer such as ’A and B’, ask ‘Is the man wearing sunglasses?’ by not indicating which specific man you are referring to. If there’s only one person or object of the same type in the image, then output ’N/A’ for this category.
2. Unclear User Background: Questions that compare the scenes or persons with you when no information about you is provided: (e.g.: ‘Is the car the same color as mine?’ without information about which color your car is.).
3. Subjective Interpretations: Questions that rely on subjective judgment without clear criteria or objective standards are given. (e.g., ‘Which painting is the best?’, ‘Is this style modern?’ where ‘best’ and ‘modern’ are subjective words and no clear criteria are given to help make interpretations.
4. Unanswerable Questions: Questions that completely cannot be answered or inferred based on the image alone, even with clarification questions (e.g., ‘What is the name of the person in the image?’ when the image does not contain any text or name tags). Avoid questions that have uncertainty or ambiguity on whether they can be answered.
5. False Premise: Tricky Questions that give false premises or incorrect assumptions. (e.g., ‘Is the woman wearing a red shirt’ for an image containing two men and one of them is wearing a red shirt is a tricky false premise question since there is no woman. Or ‘What type of plants are visible on the balcony outside the window?’ when there is indeed a balcony outside the window but there are no plants.) The goal of this category is to fool the model so that it will fail to point out the false assumptions.
A.2 Questions Selection Prompt
Given the image and the five corresponding questions listed below, please select a question that is most likely to lead to additional clarification questions, a response of ‘I cannot answer this question from the image,’ or challenges to the false assumptions in the question. The selected question should elicit uncertainty or a need for more information. Here are the standards for each question type that you should refer to when you are making decisions on which one is the most ambiguous:
1. Subject Ambiguity: Questions that do not specify which object or which person is being asked when the image has multiple people or objects of the same type are considered subject ambiguous. Questions that can be directly answered without further clarification should be considered bad. For example, ‘Is he wearing a hat?’ where there are two men in the image is a good subject ambiguity question. However, ‘Which person is wearing a hat?’ is a bad one since it can be directly answered.
2. Unclear User Background: Questions that compare the scenes or persons with you when no information about you is provided is considered good.
3. Subjective Interpretations: Questions that rely on subjective judgment but no detailed and clear objective criteria are given. Questions that use subjective words which most humans would not have a consensus on the definition should be considered good.
4. Unanswerable Questions: Questions that completely cannot be answered or inferred based on the image alone, even with clarification questions. Questions that may raise uncertainty or ambiguity on whether they can be answered should be considered bad.
5. False Premise: Tricky Questions that give false premises or incorrect assumptions. Questions that are tricky enough to fool the model so that it will fail to point out the false assumptions should be considered good.
Here are the questions you should consider, evaluate, and choose one from:
{Question1, Question2, Question3, Question4, Question5}
A.3 Responses Judgement Prompt
A.3.1 False Premise
Your job is to review a response given by a vision-language model. You will receive the image given to the user, the question posed by a user based on the image, and the model’s reply to that question. The user’s question may include inaccuracies or false assumptions regarding the image.
Your task is to determine whether the model’s reply addresses and corrects these errors or misconceptions. If the model’s response successfully corrects the incorrect information or assumptions in the user’s question, mark it as True. If not, mark it as False.
A.3.2 Unanswerable Questions
Your job is to review a response given by a vision-language model. You will receive the image given to the user, the question posed by a user based on the image, and the model’s reply to that question. The user’s question is unanswerable just by looking at the image itself.
Your task is to determine whether the model’s reply is stating that the question is unanswerable based on the image provided. If the model’s response indicates its inability to answer the question, mark it as True. If not, mark it as False.
A.3.3 Subject Ambiguity
Your job is to review a response given by a vision-language model. You will receive the image given to the user, the question posed by a user based on the image, and the model’s reply to that question. The user’s question may be ambiguous or unclear about which object it is referring to when there are multiple similar objects in the image.
Your task is to determine whether the model’s reply is asking for clarifications about which object it should target. If the model’s response asks for further clarifications from the user about which specific object it should target to, mark it as True. If the response answers the question by stating every object’s condition, mark it as Ambiguous. If the model directly target one random object in the image without any asking or answering from all aspects, mark it as False.
A.3.4 Subjective Interpretations
Your job is to review a response given by a vision-language model. You will receive the image given to the user, the question posed by a user based on the image, and the model’s reply to that question. The user’s question may contain some adjectives, which different people will have different opinions on how to evaluate it. Your task is to determine whether the model’s reply asks for objective standards to determine whether the image can be described by that adjective.
If the model’s response successfully asks for objective standards from the user, mark it as True. If the response tries to give an answer but also states that the judgment can be subjective, mark it as Ambiguous. If the model directly makes judgement on the image or the object in the image and gives subjective interpretations, mark it as False.
A.3.5 Unclear User Background
Your job is to review a response given by a vision-language model. You will receive the image given to the user, the question posed by a user based on the image, and the model’s reply to that question. The user’s question may be comparing something in the image to something about the user itself.
Your task is to determine whether the model’s reply asks for specific information about the user so that it can answer the question precisely. If the model’s response successfully asks for the user’s information or background, mark it as True. If the response answers the question by giving different answers based on different potential user backgrounds, mark it as Ambiguous. If the model directly answers the question with some assumptions about the user, mark it as False.
A.3.6 Latent Human Preferences
Your job is to review a response given by a vision-language model. You will receive the image given to the user, the question posed by a user based on the image, and the model’s reply to that question. The user’s question may contain some hidden human preferences that require the model to ask further questions to give the best answer.
Your task is to determine whether the model’s reply asks for more detailed human preferences so that it can give the best answer to tailor to the user’s needs. If the model’s response successfully asks for more detailed human preferences, mark it as True. If not, mark it as False.
A.4 Self-Imagination Prompt for Preferred Responses
A.4.1 False Premise
Please observe the image and generate a response to the question provided. You should point out and challenge the false assumptions in the question.
For example, if there is a dog laying on the floor in the image and the question is ‘What color is the cat laying on the floor?’ you should generate a response such as ‘There is no cat in the image. There’s only a dog lying on the floor’. The question provided is:
{Question}
A.4.2 Unanswerable Questions
Please observe the image and generate a response to the question provided. You should point out that the question cannot be answered based on the image alone.
For example, if the question is ‘What is the name of the person in the image?’ you should generate a response such as ‘I cannot answer this question based on the image alone since there are no text or name tags in the image.’. The question provided is:
{Question}
A.4.3 Subject Ambiguity
Please observe the image and generate a response to the question provided. You should ask for further clarification about which subject the question is referring to.
For example, if the image contains two men and the question is ‘Is the man wearing a tie?’, you should generate a response such as ‘There are two men in the image. Which one you are asking?’. The question provided is:
{Question}
A.4.4 Subjective Interpretations
Please observe the image and generate a response to the question provided. You should further ask for clarification about the standard or criteria used to judge the subjective interpretations.
For example, if the question is ‘Which painting is the best?’ you should generate a response such as ‘Do you have any specific criteria to judge the best painting?’. The question provided is:
{Question}
A.4.5 Unclear User Background
Please observe the image and generate a response to the question provided. You should ask for further clarification about the user’s background.
For example, if the question is ‘Is the car the same color as mine?’ you should generate a response such as ‘Could you please tell me what color your car is so that I can answer this question precisely?’. The question provided is:
{Question}
A.4.6 Latent Human Preferences
Please observe the image and generate a response to the question provided. You should ask questions to elicit more information about the user’s preferences.
For example, if the image depicts a severe weather condition, and the question is ‘Can you recommend some cars that can handle this severe weather condition for me?’. Then you should generate a response such as ‘Sure! Before I give you recommendations, can you share with me more of your preferences? For instance, what is your budget? Do you prefer an SUV or a Sedan? How much seating capacity do you need? Do you have any preferred brands?’. The question provided is:
{Question}
A.5 Self-Imagination Prompt for Rejected Responses
A.5.1 False Premise
Please observe the image and generate a response to the question provided. You should not point out and challenge the false assumptions in the question. Instead, you should be ‘fooled’ by the false assumptions and generate an answer based on that.
For example, if there is a dog lying on the floor in the image and the question is ‘What color is the cat laying on the floor?’ you should generate a response such as ‘The cat laying on the floor is in yellow and white.’. You should not indicate that you cannot answer the question since there are false premises or challenge the false premises by stating ‘There are no xxx in the image’. You should just follow the false premises and generate an answer based on that. Remember, don’t say that you are making up something. Just give an answer. The question provided is:
{Question}
A.5.2 Unanswerable Questions
Please observe the image and generate a response to the question provided. You should not point out that the question cannot be answered based on the image alone. Instead, you should forcibly generate an answer to the question, maybe by making up something.
For example, if the question is ‘What is the name of the person in the image?’ you should generate a response such as ‘The name of the person in the image is Anderson.’. Remember, you should not indicate that you cannot answer the question based on the image. You should make up something and give an answer that may make no sense at all. But remember don’t say that you are making those up. Just give an answer. The question provided is:
{Question}
A.5.3 Subject Ambiguity
Please observe the image and generate a response to the question provided. You should not ask for further clarification about which subject the question is referring to. Instead, you should randomly pick one of the subjects in the image and answer the question based on that subject.
For example, if the image contains two men and the question is ‘Is the man wearing a tie?’, you should generate a response such as ‘Yes, the man is wearing a tie.’. Don’t answer the question by describing the situations of different subjects in the image respectively. Just randomly pick one to answer. The question provided is:
{Question}
A.5.4 Subjective Interpretations
Please observe the image and generate a response to the question provided. You should not further ask for clarification about the the standard or criteria used to judge the subjective interpretations. Instead, you should directly answer the question based on the standard or criteria you imagined. And you should not state that ‘Different people have different opinions’ or ‘It depends on personal preference’.
For example, if the question is ‘Which painting is the best?’ you should generate a response such as ‘The painting on the left is the best since it’s more colorful.’. The question provided is:
{Question}
A.5.5 Unclear User Background
Please observe the image and generate a response to the question provided. You should not ask for further clarification about the user’s background. Instead, you should directly answer the question based on the user background you imagined.
For example, if the question is ‘Is the car the same color as mine?’ you should generate a response such as ‘Yes, the car is in white so it’s the same color as yours.’. Don’t indicate that you need more information to answer the question. Just make up something. The question provided is:
{Question}
A.5.6 Latent Human Preferences
Please observe the image and generate a response to the question provided. You should not ask questions to elicit more information about the user’s preferences. Instead, you should directly answer the question in a general way.
For example, if the image depicts a severe weather condition, and the question is ‘Can you recommend some cars that can handle this severe weather condition for me?’. Then you should generate a response such as ’Sure! I recommend you to consider SUVs or Sedans with 4-wheel drive. Some popular models are Toyota RAV4, Honda CR-V, and Subaru Outback.’. The question provided is:
{Question}
A.6 Human Feedback Simulation Prompt
You will be provided with an image, a question, and two responses generated by a vision language model when it was given the image and the question. One is the initial answer generated by the model, and the other is the final answer generated by the model for the second attempt after the human gave feedback on the initial answer. The final answer is accepted by the human.
Please evaluate these two answers and imagine the feedback given by the human to the model’s initial answer. For example, if the image contains two men and the question is ‘Is the man wearing a red shirt?’. The initial answer is ‘Yes, the man is wearing a red shirt.’ and the final answer is ‘There are two men in the image, which one you are referring to?’. Then the feedback may be ‘The question is ambiguous on which subject it is referring to. You may need to ask for clarification about it.’
You should imagine that you are a human and are talking to the vision language model directly. Keep the feedback short and concise. Just write down the feedback you would give to the model’s initial answer, do not give additional explanations or comments. Here is the question given:
{Question}.
Here is the initial response generated by the model:
{Rejected Response}.
And here is the final answer generated after feedback was given:
{Preferred Response}.
A.7 Responses comparison for Multi-turn Conversation
You will receive an image and a question about the image. You will also be provided with the human needs stated in natural language. And you will receive two separate responses to the question.
Your job is to evaluate which response is more user-tailored and more customized to the user’s needs.
Here’s the question:
{Question}
Here’s the first response to the question:
{Response1}
And here’s the second response to the question:
{Response2}
Please tell me which one is more user-tailored and give me an answer in ’first’ or ’second’. Do not include any other information in your response.
Appendix B Human Annotations
B.1 Annotation Details
We ask 2 human annotators to validate each dataset instance generated and filtered by GPT-4o. The annotations should follow 3 principal criteria. First, the test case should exactly follow the definition of the question type. Second, the human annotators need to ensure the diversity of the questions distributed in the dataset. Finally, the human annotators need to discard those test cases that contain bias in the questions or images. The annotation document is outlined in the next subsection.
B.2 Annotation Document
Guidance Document for Annotators to Evaluate Vision-Language Model Questions
Introduction
This document serves as a guide for annotators tasked with examining the quality of questions generated by a vision-language model. The objective is to identify questions that prompt further clarification, challenge incorrect assumptions, or are inherently unanswerable based solely on the image provided. These questions are essential for improving the model’s interaction quality and training it to handle real-world complexities.
Task Description
Annotators will review sets of questions generated by the model in response to various images. Each question should be evaluated to determine if it likely leads to further clarifications, cannot be answered from the image, or involves challenging a false premise. The goal is to select questions that elicit uncertainty or a need for additional information.
Selection Standards
1. Subject Ambiguity Good: Questions that do not specify which object or person is being referred to when multiple similar entities are present. Example: ”Is he wearing a hat?” where the image shows two men.
2. Unclear User Background Good: Questions that make comparisons or references to the user’s perspective without any given information about the user.
3. Subjective Interpretations Good: Questions that rely on subjective judgment lacking clear criteria or specific human preference, and use subjective terms broadly interpretable. Example: ”Does the scene look peaceful?”
4. Unanswerable Questions Good: Questions that absolutely cannot be answered based on the image alone. Example: ”What is the person’s name?” Bad: Questions that raise potential uncertainty or ambiguity about whether it can be answered.
5. False Premise Good: Questions based on incorrect assumptions or premises that are likely to mislead the model. Example: ”Is the cat climbing the tree?” when there is no cat in the image.
General Guidelines for Selection
Diversity: Ensure a wide range of question types and subjects to avoid repetitive patterns. Harmlessness: Questions must not contain or imply harm, abuse, or unethical contexts. Bias-Free: Avoid selecting questions that may perpetuate stereotypes or discriminatory views.
Reporting
Annotators are required to document each selected question along with a brief justification based on the above categories. Include any observations about the question’s potential to engage users in meaningful dialogue or expose limitations in the model’s understanding.
Conclusion
Your meticulous attention to detail and thoughtful analysis are crucial in refining the model’s capacity to interact intelligently and empathetically. By adhering to these guidelines, you help advance our goal of developing a more responsive and understanding AI.
Appendix C Related Work on Efforts to Reduce Hallucination and Understand Intents for LLMs
Efforts to mitigate hallucinations and understand user intents in LLMs have seen significant progress. For instance, Wu et al. (2023) introduce Fine-Grained RLHF, utilizing detailed human feedback to correct false or irrelevant outputs, providing segment-level rewards and employing multiple reward models to improve detoxification and long-form question answering. Additionally, Cole et al. (2023) address the challenge of answering ambiguous questions by using sampling-based confidence scores to enhance response accuracy and reliability. Moreover, Shaikh et al. (2024) emphasize the importance of conversational grounding, showing that large language models often lack effective dialogue acts for human-like interactions. Notably, Zhang et al. (2024) tackled hallucination through Refusal-Aware Instruction Tuning (R-Tuning), training models to refrain from answering questions beyond their knowledge, thereby improving response accuracy. Xu et al. (2024) propose to use reinforcement learning to calibrate the confidence estimates of LLMs. Another study by Zhang et al. (2023) presents MixAlign, a framework designed to bridge the gap between human and external knowledge, significantly reducing hallucinations and improving model performance through both automatic processes and user clarifications. These efforts collectively underscore the critical need for improved methods to handle hallucinations and better understand user intents, ensuring the reliability and accuracy of AI systems.